In simultaneous equation, uing substitutional method, solve 3x+3y=10 and 2x-y=2
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

In simultaneous equation, using substitution method, solve
3x + 3y = 10
2x - y = 2
:
Rearrange the 2nd equation for substitution
subtract 2x from both sides
-y = -2x + 2
multiply by -1
y = 2x - 2
:
Replace y with (2x-2) in the 1st equation
3x + 3(2x - 2) = 10
3x + 6x - 6 = 10
9x = 10 + 6
9x = 16
x = 16/9
:
Find y
y = 2(16/9) - 2
y = (32/9) - (18/9)
y = (14/9)
:
Solution
x = 16/9
y = 14/9
:
:
:
Check this on your calc in the 1st equation
3(16/9) + 3(14/9) =

Answer by ikleyn(52787)

.
Regarding the Substitution method, Elimination method, Cramer's method for solving linear systems of algebraic equations with two unknowns see the lessons
    - Solution of the linear system of two equations in two unknowns by the Substitution method
    - Solution of the linear system of two equations in two unknowns by the Elimination method
    - Geometric interpretation of the linear system of two equations in two unknowns
    - Solving word problems using linear systems of two equations in two unknowns

    - Determinant of a 2x2-matrix
    - Solving systems of linear equations in two unknowns using the Cramer's rule
    - Solving word problems by the Cramer's rule after reducing to systems of linear equations in two unknowns
    - OVERVIEW of LESSONS on determinants of 2x2-matrices and Cramer's rule for systems in 2 unknowns
in this site.